Project

General

Profile

NDOS Commissioning Meeting 2011/04/07

Attendance: Jon, Andrew, Jarek, Minerba, Enrique, Mark Messier, Gary, Alena, Athans, Leon, Jan, Nick, Gavin, Susan, Tricia

Agenda:

  1. NDOS update
  2. DAQ/Electronics Update & Plans
  3. DCS Status & Plans
  4. DataCheck update
  5. DSO Studies
  6. AOB

NDOS Status:

  • Mark: APD installation completed on top. Three APDs will be installed on the side today (now). APD installation will continue on the side after top-off is complete.
  • Denis will modify code for DSO scans and threshold calculations; will run DSO scan in the morning.

DAQ/Electronics Status & Plans:

  • Leon: found a problem on one of the DCM timing links. Chip was fried. Looking into the cause, but most likely due to an ethernet cable being plugged in (same physical form). DCM was replaced (dcm-06 -> dcm-03). Denis commented that it was relatively painless to make the change in the software.
  • Peter:
    • upgrade to new version of DDS has been deemed a success, so the software has been merged and put into use.
    • Kurt has a new application manager that will be used in the CR. This allows one to reboot DCMs
    • Currently struggling with spill server, Holger is working on it.
    • Ron R. wants to test new kernel on DCMs; will likely wait until next week.

DCS Status & Plans:

  • Chad reported via email: "... This morning I was out at the NDOS and I moved the computer from rack 1 to rack 3 (the middle rack). I also installed a patch panel (no wires connected to it yet) in and metal plates over the space left behind by the computer in rack 1."
  • Athans: finalized DCM monitor code but is having problems running his code lately. Discussion about DAQ/DCS interface ensued. Decided to continue this discussion at next week's collaboration meeting.

DataCheck Update (doc-db 5795)

AOB

  • Mark reported that Chad has found a timing peak in the Booster data around 332-334 us. Results will be shown at tomorrow's offline meeting. Hurray! Kudos to Chad, Minerba and Denis.